Find the general term aVn of a sequence whose first four terms are giving: 5,9,13,17,...for n=1,2,3,4,... Please help me with this problem.
Answer by venugopalramana(3286)

Find the general term aVn of a sequence whose first four terms are giving: 5,9,13,17,...for n=1,2,3,4,... Please help me with this problem.
WE FIND
9-5=13-9=17-13=4...SO THIS IS A.P
A=5 AND D=4
GENERAL TERM
TN=A+(N-1)D=5+(N-1)4=5+4N-4=4N+1